OpenStudy (anonymous):

A recursive function is shown below: f(1) = 4 and f(n) = f(n - 1) - 8; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 4, -4, -12, -20, -28, ... 4, 4, 12, 20, 28, ... 4, 12, 20, 28, 36, ... 4, -12, -20, -28, -36, ...

OpenStudy (anonymous):

@cj49 ?

OpenStudy (cj49):

f(2)=f(1)=-8=4-8=-4